OpenStudy (anonymous):

find the maximum y-value on the graph of y=f(x) f(x)=-x^2+2x+5

OpenStudy (anonymous):

can u find the vertex of this parabola?

OpenStudy (leksi):

f'=-2x+2 f'=0 -2x+2=0 x=1 max =(-(-1))^2+2*(-1)+5=4

OpenStudy (anonymous):

so the answer is 4

OpenStudy (leksi):

yes

OpenStudy (anonymous):

thanks so much

OpenStudy (leksi):

bye